# Watch Your Step: Optimal Retrieval for Continual Learning at Scale

### v1

@@ Abstract
One of the most widely used approaches in continual learning is referred to as replay. Replay methods support interleaved learning by storing past experiences in a replay buffer. Although there are methods for selectively constructing the buffer and reprocessing its contents, there is limited exploration of the problem of selectively retrieving samples from the buffer. Current solutions have been tested in limited settings and, more importantly, in isolation. Existing work has also not explored the impact of duplicate replays on performance. In this work, we propose a framework for evaluating selective retrieval strategies, categorized by simple, independent class- and sample-selective primitives. We evaluated several combinations of existing strategies for selective retrieval and present their performances. Furthermore, we propose a set of strategies to prevent duplicate replays and explore whether new samples with low loss values can be learned without replay. In an effort to match our problem setting to a realistic continual learning pipeline, we restrict our experiments to a setting involving a large, pre-trained, open vocabulary object detection model, which is fully fine-tuned on a sequence of 15 datasets.
@@ 1 Introduction
Motivated by recent work showing abrupt representation drift upon entering a new task [35, 8], we then experiment with the idea of only using replay for a new datapoint when its loss is above a certain threshold. Here, we find that maintaining at least a 1:1 ratio between new samples and replay samples is critical for overall performance, as even witholding replay samples for new samples in the bottom 3% of the loss magnitudes leads to significant drops in performance.
@@ 2 Related Work
Figure 1: The problem settings of continual learning. Top: the original and most common setup, where dataset is divided into N balanced subsets and the model is trained sequentially on each subset. Middle: another common setup, which is the same as above except the model is pre-trained; forgetting on pre-training tasks is ignored. Bottom: our setup, where a pre-trained model is sequentially trained on OOD datasets and forgetting on pre-training tasks is minimized.
@@ 4.2 Sample Selection Primitives
Adversarial Shapley Value [39]: Each sample in a selected candidate set is scored according to the Adversarial Shapley Value (ASV), which is derived from the KNN Shapley Value (KNN-SV) [21]. The KNN Shapley Value is an efficient approximation of the Shapley Value [38] used in data valuation and can be understood as an algorithto attributeng predictions made by a model to individual samples in an evaluation set (e.g. a validation set or, in this case, a batch of new data) to individual samples in a training set (which, in our case, is instead the candidate set drawn from our replay buffer). When computing a set of KNN-SVs for a candidate set $D_{c}$ with $N_{c}$ images and an evaluation set $D_{e}$ with $N_{e}$ images, our goal is to compute a $(N_{c},N_{e})$ tensor of KNN-SVs, where a high KNN-SV at index $(i,j)$ means that the model’s prediction for the evaluation image $j$ is highly attributable to candidate image $i$ being present in the training dataset.
@@ 4.5 Loss Adaptivity
New samples with low losses have smaller gradients and thus may be able to be learned without replay. Limiting the replay to new samples with high loss values can reduce the number of forward and backward passes computed on the replay samples and potentially improve the learning of new tasks. Here, we propose an algorithm to reduce the number of replayed samples and for reducing the magnitude of replayed samples’ losses.
@@ 5.3 Preventing Duplicate Replays
We find the optimality of per-dataset deduplication to be unsurprising, as each new dataset represents a distinct continual learning problem (see Figure 6) which may call for maximum selectivity, especially in the early epochs since that is when each new sample’s gradients are largest and have the highest variance.
@@ 5.4 Loss Adaptivity
Following our main experiments, we tested whether keeping a 1:1 ratio between new and replayed samples, regardless of new samples’ losses, is necessary. As seen in Table 6, witholding replay samples for as little as 3% of new samples leads to a performance decrease.
@@ 7 Conclusion
In this work, we modify and evaluate algorithms for selective retrieval in continual learning. In addition, we develop simple, yet effective algorithms for deduplication and replay buffer construction, both of which are particularly useful in our problem setting. We then provided insights on the prospect of a loss-adaptive replay ratio, the shapes of algorithms’ distributions, and dataset-dependent forgetting.

### v2

@@ Abstract
In continual learning, a model learns incrementally over time while minimizing interference between old and new tasks. One of the most widely used approaches in continual learning is referred to as replay. Replay methods support interleaved learning by storing past experiences in a replay buffer. Although there are methods for selectively constructing the buffer and reprocessing its contents, there is limited exploration of the problem of selectively retrieving samples from the buffer. Current solutions have been tested in limited settings and, more importantly, in isolation. Existing work has also not explored the impact of duplicate replays on performance. In this work, we propose a framework for evaluating selective retrieval strategies, categorized by simple, independent class- and sample-selective primitives. We evaluated several combinations of existing strategies for selective retrieval and present their performances. Furthermore, we propose a set of strategies to prevent duplicate replays and explore whether new samples with low loss values can be learned without replay. In an effort to match our problem setting to a realistic continual learning pipeline, we restrict our experiments to a setting involving a large, pre-trained, open vocabulary object detection model, which is fully fine-tuned on a sequence of 15 datasets.
@@ 1 Introduction
Motivated by recent work showing abrupt representation drift upon entering a new task [35, 8], we then experiment with the idea of reducing the impact of replay when new samples’ losses are below a threshold, along with reducing the number of replays in these cases. This procedure reflects the hypothesis that reducing the impact of replay for new samples with the smallest gradients can lead to extra plasticity for new tasks, without incurring more forgetting of previous tasks. Here, we find that applying this procedure rapidly leads to forgetting previous tasks, as the case where only 3% of new samples are below the loss threshold leads to significant drops in performance.
@@ 2 Related Work
Figure 1: The problem settings of continual learning research. Top: the original and most common setup, where dataset is divided into N balanced subsets and the model is trained sequentially on each subset. Middle: another common setup, which is the same as above except the model is pre-trained; forgetting on pre-training tasks is ignored. Bottom: our setup, where a pre-trained model is sequentially trained on OOD datasets and forgetting on pre-training tasks is minimized.
@@ 4.2 Sample Selection Primitives
Adversarial Shapley Value [39]: Each sample in a selected candidate set is scored according to the Adversarial Shapley Value (ASV), which is derived from the KNN Shapley Value (KNN-SV) [21]. The KNN Shapley Value is an efficient approximation of the Shapley Value [38] used in data valuation and can be understood as an algorithm for attributing predictions made by a model for individual samples in an evaluation set (e.g. a validation set or, in this case, a batch of new data) to individual samples in a training set (which, in our case, is instead the candidate set drawn from our replay buffer). When computing a set of KNN-SVs for a candidate set $D_{c}$ with $N_{c}$ images and an evaluation set $D_{e}$ with $N_{e}$ images, our goal is to compute a $(N_{c},N_{e})$ tensor of KNN-SVs, where a high KNN-SV at index $(i,j)$ means that the model’s prediction for the evaluation image $j$ is highly attributable to candidate image $i$ being present in the training dataset.
@@ 4.5 Loss Adaptivity
New samples with low losses have smaller gradients and thus may be able to be learned without replay. Selectively limiting the impact of replay in this way could lead to more plasticity for new tasks, without incurring more forgetting. Here, we propose an algorithm for adaptively reducing the number of replayed samples and for reducing the magnitude of replayed samples’ losses according to each new sample’s loss magnitiude.
@@ 5.3 Preventing Duplicate Replays
This result reflects the fact that each new dataset represents a distinct continual learning problem (see Figure 6) which may call for maximum selectivity in the early epochs since that is when each new sample’s gradients are largest and have the highest variance. This selectivity can then be reduced in later epochs in favor of greater coverage of the replay buffer.
@@ 5.4 Loss Adaptivity
Following our main experiments, we tested whether keeping a 1:1 ratio between new and replayed samples and maintaining the full impact of the replay loss, regardless of new samples’ losses, is necessary. As seen in Table 4, the case where as little as 3% of new samples leads to a performance decrease.
@@ 6.1 The Shape of SWIL and GRASP
Importantly, this insight leads to the conclusion that SWIL may scale quite well with the number of classes in the replay buffer, as it would then be more likely to produce lower-entropy distributions over classes in the replay buffer. Such an effect could lead SWIL to outperform GRASP in certain scenarios, although the inverse would likely hold as well.
@@ 7 Conclusion
Selective retrieval is a key component of any replay algorithm. Existing methods can be organized into class- and sample-selective primitives; however, the best class- and sample-selective primitive cannot be combined to reach a new state of the art. This may be explained by the best selective retrieval algorithms being highly selective at times, leading to an overall insufficient amount of randomness across the learning trajectory.
Any algorithm used for selective retrieval should ensure that no duplicates are retrieved within a single continual learning task (dataset). This leads to retrieval being highly selective during the early epochs for each task (when losses on new samples are highest) and greater coverage of the replay buffer in later epochs.
When using replay, it is essential that the replay loss’s magnitude is kept consistent, regardless of new samples’ losses. This reflects the necessity of maintaining existing representations in the network, especially in the problem setting of continual learning at scale, since previous representations cover many different classes and are required for forward transfer (optimal learning of new tasks).
Extensions to this work includes measures for reducing the computational overhead associated with replaying full inputs. This can be achieved by developing methods which replay old samples on the level of tokens, with an architecture similar to that of retrieval augmented generation [26].
One key limitation of this work is that the replay buffer was not updated with samples from downstream tasks. Adding new samples to the replay buffer and ensuring they are replayed could potentially remove the performance gap between the ODinW performance upper bound and ODinW performance achieved when using replay.
Another key limitation of this work is that the continual learning phase only consisted of a sequence of 15 datasets. In practice, this sequence would need to be much longer for sequential learning to make sense, as all 15 datasets could be learned jointly. Some interesting properties of the problem setting also may only be revealed given a task sequence of sufficient length.
Based on the preceding analyses on dataset-dependent forgetting, a higher replay ratio for smaller downstream datasets should also be explored.
8 Acknowledgments
This work is partially supported by the NSF EFRI BRAID grant #2317706.